In 2008,the world population was 6.7 billion and the exponential growth rate was 1.14% per year.
A. Find the exponential growth function
B. Predict the world population in 2014
C. When will the world population be 8.0 billion?

Answer by stanbon(57361)

In 2008,the world population was 6.7 billion and the exponential growth rate was 1.14% per year.
A. Find the exponential growth function
A(t) = 6.7 bil (1.14)^t where t is the number of years after 2008.
----------------------------------
B. Predict the world population in 2014
A(6) = 6.7 bil (1.14)^6 = 6.7 bil (2.1950) = 14.71 billion
-------------------------
C. When will the world population be 8.0 billion
8 bil = 6.7 bil (1.14)^t
---
1.194 = 1.14^t
---
Take the log:
t*log(1.14) = log(1.194)
t = 1.35 years
=====================
Cheers,
Stan H.
=====================